OpenStudy (anonymous):

determine whether f(x) represents a polynomial function . If possible identify the degree and type of polynomial function f(x)=1/3-x is f(x) a polynomial function? no yes it is a binomial with degree 2 yes it is a binomial with degree 0 yes it is a binomial with degree 1 which on is right

OpenStudy (anonymous):

degree 0

OpenStudy (campbell_st):

quick question is f(x) 1. \[f(x) = \frac{1}{3} - x\] or 2. \[f(x) = \frac{1}{3 - x}\]

OpenStudy (anonymous):

oh good question

OpenStudy (anonymous):

0k it 3

OpenStudy (campbell_st):

a zero degree polynomial is a constant as in the example below \[f(x) = 3x^0 = 3(1) = 3\]

OpenStudy (anonymous):

campbell what program are you using/how do you write the formulas in chat?

OpenStudy (campbell_st):

I just use the equation option at the bottom, its a latex editor...

OpenStudy (anonymous):

f(x)=1/3-x

OpenStudy (anonymous):

oops made a typo here it is again fixed x = degree 1 x^2 = degree 2 x^3 = degree 3 x^2-2x =degree 3